Jamaica Corp. is adding a new assembly line at a cost of $8.5 million. The firm expects the project to generate cash flows of $2 million, $3 million, $4 million, and $5 million over the next four years. Its cost of capital is 16 percent. Net present value: What is the MIRR and should company add the new assembley line?

18.58%, no
18.57%, yes
18%, no
19.87%, no
Answer #1

NPV = -8.5 + 2/(1.16) + 3/(1.16)2 + 4/(1.16)3 + 5/(1.16)4

NPV = $0.78 million

MIRR = (FV of Cash inflow/PV of Cash out flow)1/n - 1

FV of Cash inflow = 2(1.16)3 + 3(1.16)2 + 4(1.16) + 5

FV of Cash inflow = $16.80 million

MIRR = (16.80/8.5)1/4 - 1

MIRR = 18.57%

Yes, investment should be made as NPV is positive and MIRR is greater than cost of capital